# Vision-Based Mistake Analysis in Procedural Activities: A Review of Advances and Challenges

### v1

@@ VII-C3 Unsupervised Methods
Existing unsupervised methods follow strategies centered on predictive modeling, where states of modalities (e.g., gaze, human 2D/3D pose) are predicted and mistakes inferred from large prediction deviations. They rely on latent modeling of correct trajectories, often as feature embeddings, where deviations from learned distributions are flagged as mistakes/anomalies. This category of mistake recognition/detection methods is related to VAD methods (see Section II-A). However, even though they do not explicitly model task semantics, action steps, or procedural knowledge, unsupervised mistake analysis methods are applied to procedural activities (e.g., cooking, object assembly), where mistakes are subtle and often arise from missing, improperly ordered, or poorly executed steps. In contrast, VAD methods with similar approaches [68, 69, 70, 71] focus on coarse distinct actions with no procedural context, usually in surveillance applications. Mazzamuto et al. [46] leverages egocentric gaze prediction as an unsupervised proxy for task regularity. It defines a novel gaze completion task and measures deviations between predicted and observed gaze to detect mistakes online, without using any labels or manual annotations. To capture this, they formulate the gaze completion task with a transformer-based model that learns to predict future gaze trajectories based on past egocentric video and gaze history. During inference, discrepancies between the predicted gaze and the actual observed gaze are interpreted as anomalies, with large deviations signaling possible mistakes or disruptions in task execution.
@@ VIII-A Challenges in Mistake Analysis
Ambiguity over permissible variations & actual errors: A key problem in procedural mistake analysis is distinguishing legitimate execution variations from true errors. Human procedures are inherently variable, often allowing multiple valid action sequences, interaction styles, or tool-use alternatives achieving the same goal without compromising quality. Annotation challenges stem from limited consensus on what constitutes a mistake versus a permissible variation, driving high labeling costs and inter-annotator disagreement, particularly when procedural flexibility is high or correctness criteria implicit. Consequently, existing datasets often omit such nuance or adopt simplified taxonomies overlooking real-world task complexity. Modeling approaches typically rely on distributional regularities learned from data, lacking explicit representations of task goals, procedural constraints, or causal dependencies. This leads to misclassification of rare but correct behaviors and missed detections of context-dependent violations requiring deeper task reasoning. Addressing this necessitates models to reason about action affordances, object roles, and the causal effects of omissions or order errors. Recent video-language models [72] jointly encode temporal and semantic context and can be prompted with structured task goals or contextual cues. Yet precise error attribution remains difficult, especially for finer distinctions and task-specific correctness rules. Vision–language approaches like PREGO [24] and TI-PREGO [37] show promise in identifying procedural errors, though executional ones remain underexplored. Their sensitivity to prompt phrasing and contextual formulation [73] further raises robustness concerns in real-world settings.
@@
Disjoint embedding spaces between LLMs & video encoders: A key limitation in current mistake analysis frameworks lies in the lack of task-level reasoning, where models fail to encode procedural goals, causal dependencies, and object affordances, resulting in systems that primarily detect low-level visual deviations without understanding whether an observed deviation truly constitutes a mistake [46]. Video-Language Models (Video LLMs) [24, 37] have emerged as a promising direction, integrating the semantic richness of LLMs with the temporal and spatial modeling capacity of video encoders. However, representational misalignment between the two modalities (vision, language) remains a core challenge. Current designs rely on dual-stream architectures with shallow fusion, handcrafted prompts, or minimal supervision of cross-modal correspondences. As a result, they struggle to ground complex procedural logic or contextual correctness in visual data [74] and remain limited to detecting procedural mistakes, which require alignment with predefined step sequences, rather than executional mistakes, which require fine-grained perception of action quality and context-dependent deviations.
@@ VIII-B Future Directions in Mistake Analysis
Lifelong learning & open-vocabulary recognition: A major limitation of current methods is the reliance on closed-world training setups, where action-object classes, and task flows are assumed fixed and fully known during training. However, real-world procedures are dynamic, evolving across domains, involving diverse object-tool configurations, and introducing novel behaviors or error patterns unseen during training. This motivates a shift toward lifelong learning and open-vocabulary recognition frameworks for more robust, scalable, and generalizable analysis. Lifelong learning paradigms, particularly for video understanding (e.g., [75]), enable models to incrementally learn from streaming video while retaining performance on prior tasks. Such adaptability is crucial for vision-based mistake analysis, especially for executional mistakes, where new action styles may arise due to tool substitutions. Complementing continual learning, open-vocabulary recognition allows models to describe and detect mistakes beyond closed taxonomies, enabling zero-shot recognition of novel mistake classes. Future methods could draw inspiration from action understanding works that generalize beyond fixed label sets using vision-language alignment and compositional prompts in multi-modal large language models (e.g., [76, 77]).
@@
Explainable analysis via symbolic reasoning integration with video models: A key challenge in vision-based mistake analysis is the internal mechanisms that guide model decisions, especially when detecting subtle deviations in procedural tasks. Current end-to-end neural models often operate as black boxes, offering little insight into why a given action is classified as a mistake. Such explainability should be considered a foundational principle and a critical requirement due to the inherent flexibility of human-performed tasks. Many procedures allow for multiple valid execution paths, making it essential for models to differentiate between permissible variations and true violations of task constraints. A promising direction towards this goal is integrating symbolic reasoning, such as procedural constraints, causal rules and temporal logic, with neural video perception. Mistake analysis could benefit from neuro-symbolic designs that (a) parse observed video into symbolic representations of steps (a concept briefly explored in PREGO [24] and TI-PREGO [37] for aligning observed actions with expected procedural steps) and object states [78], (b) enforce procedural logic via symbolic reasoners or temporal logic automata [79], and (c) ground discrepancies (both procedural and executional) as violations of explicit task rules. Such hybrid systems would enhance detection accuracy and offer interpretable explanations as well as zero-shot generalization across new procedures.
@@
Understanding error dependencies & propagation:Building on interpretable representations for procedural understanding, the next step is to move beyond detection and examine how errors influence downstream actions within task sequences. In real-world industrial or medical settings, errors rarely occur in isolation; initial deviations can propagate, altering execution and causing more severe failures. Understanding these inter-error dependencies is essential for systems that not only detect mistakes but anticipate and mitigate cascading effects. Existing approaches largely focus on individual errors, often overlooking their influence on subsequent ones. Future work should model such dependencies using causal inference and probabilistic graphical models. By learning the likelihood of specific error transitions (e.g., Error A increases the probability of Error C), these methods could enable early interventions to prevent cascading failures. Integrating predictive capabilities into real-time mistake detection could enhance automated quality control and prevention. Recent synthetic-domain works like [80], can model explicit visual state transitions, providing a foundation for inferring causal chains of mistakes via object state transitions in real-world procedural tasks.
@@
Enhancing datasets with counterfactual state transitions and synthetic mistakes: Procedure-aware video understanding and mistake recognition require modeling both actual and hypothetical scene state transitions, capturing action-induced transformations and plausible deviations from incomplete or erroneous actions. Several works have highlighted the link between action or activity recognition and object/scene state changes [81, 82, 83, 84, 85]. A promising direction for advancing mistake analysis involves augmenting procedural datasets with synthetic state-change sequences and counterfactual exemplars [26, 86, 87, 88], enabling models to reason about the causal and semantic impact of actions by grounding visual representations in expected and unexpected outcomes. This fosters richer temporal representations that distinguish correct progressions from deviations or failures, even in subtle or ambiguous cases. Recent work by Kung et al. [26] showed that, given prior knowledge of an activity’s execution protocol, one can use structured procedural knowledge bases (e.g., WikiHow [89]) or LLM-generated textual descriptions to define expected pre- and post-action states and hypothetical post-conditions for potential failures. These textual representations can then be aligned with visual features to reinforce plausible state transitions while decoupling them from counterfactual ones via contrastive learning, thereby strengthening the visual grounding of procedural semantics.
@@ IX Conclusions
Vision-based mistake analysis in procedural activities is a rapidly evolving field with potential to improve task performance, safety, and efficiency across domains. This paper reviews existing methods, datasets, and evaluation metrics, outlining challenges and opportunities in detecting and predicting procedural and executional errors. By categorizing approaches by procedural structure, supervision, and learning strategies, we offer a structured view of the state of the art. Key challenges remain—ambiguity between permissible variations and true mistakes, limited comprehensive datasets, and the need for models that reason about error dependencies and propagation. Addressing these requires innovations such as neuro-symbolic reasoning, leveraging large language models for procedural understanding, and enriching datasets with counterfactual transitions and synthetic mistakes. Future research should emphasize explainable mistake analysis, stronger cross-task generalization, and real-time predictive detection, unlocking new possibilities for human–robot collaboration, automation, and assistive technologies, and ultimately transforming how mistakes are understood and mitigated in structured activities.

### v2

@@ VII-C3 Unsupervised Methods
Existing unsupervised methods follow strategies centered on predictive modeling, where states of modalities (e.g., gaze, human 2D/3D pose) are predicted and mistakes inferred from large prediction deviations. They rely on latent modeling of correct trajectories, often as feature embeddings, where deviations from learned distributions are flagged as mistakes/anomalies. This category of mistake recognition/detection methods is related to VAD methods (see Section II-A). However, even though they do not explicitly model task semantics, action steps, or procedural knowledge, unsupervised mistake analysis methods are applied to procedural activities (e.g., cooking, object assembly), where mistakes are subtle and often arise from missing, improperly ordered, or poorly executed steps. In contrast, VAD methods with similar approaches [sato2023prompt, markovitz2020graph, noghre2024exploratory, stergiou2024holistic] focus on coarse distinct actions with no procedural context, usually in surveillance applications. Mazzamuto et al. [mazzamuto2025gazing] leverages egocentric gaze prediction as an unsupervised proxy for task regularity. It defines a novel gaze completion task and measures deviations between predicted and observed gaze to detect mistakes online, without using any labels or manual annotations. To capture this, they formulate the gaze completion task with a transformer-based model that learns to predict future gaze trajectories based on past egocentric video and gaze history. During inference, discrepancies between the predicted gaze and the actual observed gaze are interpreted as anomalies, with large deviations signaling possible mistakes or disruptions in task execution.
@@ VIII-A Challenges in Mistake Analysis
Ambiguity over permissible variations & actual errors: A key problem in procedural mistake analysis is distinguishing legitimate execution variations from true errors. Human procedures are inherently variable, often allowing multiple valid action sequences, interaction styles, or tool-use alternatives achieving the same goal without compromising quality. Annotation challenges stem from limited consensus on what constitutes a mistake versus a permissible variation, driving high labeling costs and inter-annotator disagreement, particularly when procedural flexibility is high or correctness criteria implicit. Consequently, existing datasets often omit such nuance or adopt simplified taxonomies overlooking real-world task complexity. Modeling approaches typically rely on distributional regularities learned from data, lacking explicit representations of task goals, procedural constraints, or causal dependencies. This leads to misclassification of rare but correct behaviors and missed detections of context-dependent violations requiring deeper task reasoning. Addressing this necessitates models to reason about action affordances, object roles, and the causal effects of omissions or order errors. Recent video-language models [tang2025video] jointly encode temporal and semantic context and can be prompted with structured task goals or contextual cues. Yet precise error attribution remains difficult, especially for finer distinctions and task-specific correctness rules. Vision–language approaches like PREGO [flaborea2024prego] and TI-PREGO [plini2024ti] show promise in identifying procedural errors, though executional ones remain underexplored. Their sensitivity to prompt phrasing and contextual formulation [zhou2022learning] further raises robustness concerns in real-world settings.
@@
Disjoint embedding spaces between LLMs & video encoders: A key limitation in current mistake analysis frameworks lies in the lack of task-level reasoning, where models fail to encode procedural goals, causal dependencies, and object affordances, resulting in systems that primarily detect low-level visual deviations without understanding whether an observed deviation truly constitutes a mistake [mazzamuto2025gazing]. Video-Language Models (Video LLMs) [flaborea2024prego, plini2024ti] have emerged as a promising direction, integrating the semantic richness of LLMs with the temporal and spatial modeling capacity of video encoders. However, representational misalignment between the two modalities (vision, language) remains a core challenge. Current designs rely on dual-stream architectures with shallow fusion, handcrafted prompts, or minimal supervision of cross-modal correspondences. As a result, they struggle to ground complex procedural logic or contextual correctness in visual data [rawal2023dissecting] and remain limited to detecting procedural mistakes, which require alignment with predefined step sequences, rather than executional mistakes, which require fine-grained perception of action quality and context-dependent deviations.
@@ VIII-B Future Directions in Mistake Analysis
Lifelong learning & open-vocabulary recognition: A major limitation of current methods is the reliance on closed-world training setups, where action-object classes, and task flows are assumed fixed and fully known during training. However, real-world procedures are dynamic, evolving across domains, involving diverse object-tool configurations, and introducing novel behaviors or error patterns unseen during training. This motivates a shift toward lifelong learning and open-vocabulary recognition frameworks for more robust, scalable, and generalizable analysis. Lifelong learning paradigms, particularly for video understanding (e.g., [gowda2024continual]), enable models to incrementally learn from streaming video while retaining performance on prior tasks. Such adaptability is crucial for vision-based mistake analysis, especially for executional mistakes, where new action styles may arise due to tool substitutions. Complementing continual learning, open-vocabulary recognition allows models to describe and detect mistakes beyond closed taxonomies, enabling zero-shot recognition of novel mistake classes. Future methods could draw inspiration from action understanding works that generalize beyond fixed label sets using vision-language alignment and compositional prompts in multi-modal large language models (e.g., [gupta2024open, chatterjee2023opening]).
@@
Explainable analysis via symbolic reasoning integration with video models: A key challenge in vision-based mistake analysis is the internal mechanisms that guide model decisions, especially when detecting subtle deviations in procedural tasks. Current end-to-end neural models often operate as black boxes, offering little insight into why a given action is classified as a mistake. Such explainability should be considered a foundational principle and a critical requirement due to the inherent flexibility of human-performed tasks. Many procedures allow for multiple valid execution paths, making it essential for models to differentiate between permissible variations and true violations of task constraints. A promising direction towards this goal is integrating symbolic reasoning, such as procedural constraints, causal rules and temporal logic, with neural video perception. Mistake analysis could benefit from neuro-symbolic designs that (a) parse observed video into symbolic representations of steps (a concept briefly explored in PREGO [flaborea2024prego] and TI-PREGO [plini2024ti] for aligning observed actions with expected procedural steps) and object states [gouidis2024enabling], (b) enforce procedural logic via symbolic reasoners or temporal logic automata [choi2024towards], and (c) ground discrepancies (both procedural and executional) as violations of explicit task rules or specific semantic roles (e.g., predicate or object mismatches [li2025mistake]). Such hybrid systems would enhance detection accuracy and offer interpretable explanations as well as zero-shot generalization across new procedures. Regarding explainability, [patsch2025mistsense] demonstrate the potential of using LLMs to generate intuitive natural language error descriptions directly from visual features; however, ensuring these generative explanations remain factually grounded and structurally verifiable, remains a critical challenge that necessitates the integration of explicit symbolic constraints. Furthermore, the reliance on standard captioning metrics (e.g., BLEU, CIDEr) to evaluate these systems [patsch2025mistsense] is arguably inadequate for mistake analysis; such metrics quantify textual overlap rather than the causal correctness or interventional utility of the diagnosis, underscoring the need for new, logic-grounded evaluation protocols.
@@
Understanding error dependencies & propagation: Building on interpretable representations for procedural understanding, the next step is to move beyond detection and examine how errors influence downstream actions within task sequences. In real-world industrial or medical settings, errors rarely occur in isolation; initial deviations can propagate, altering execution and causing more severe failures. Understanding these inter-error dependencies is essential for systems that not only detect mistakes but anticipate and mitigate cascading effects. While early works like AMNAR [huang2025modeling] use graph structures to predict valid next-steps based on current history, future methods must extend this to explicitly learn the likelihood of specific error transitions (e.g., Error A increases the probability of Error C) to enable early interventions preventing cascading failures. Integrating predictive capabilities into real-time mistake detection could enhance automated quality control and prevention. Recent synthetic-domain works like [hong2021transformation], can model explicit visual state transitions, providing a foundation for inferring causal chains of mistakes via object state transitions in real-world procedural tasks.
@@
Enhancing datasets with counterfactual state transitions and synthetic mistakes: Procedure-aware video understanding and mistake recognition require modeling both actual and hypothetical scene state transitions, capturing action-induced transformations and plausible deviations from incomplete or erroneous actions. Several works have highlighted the link between action or activity recognition and object/scene state changes [ChopANDLearn, wang2016actions, Bacharidis_2023_ICCV, souvcek2022look, niu2024schema]. A promising direction for advancing mistake analysis involves augmenting procedural datasets with synthetic state-change sequences and counterfactual exemplars [kung2025changed, GenHowTo, chatzi2024counterfactual, souvcek2025showhowto], enabling models to reason about the causal and semantic impact of actions by grounding visual representations in expected and unexpected outcomes. This fosters richer temporal representations that distinguish correct progressions from deviations or failures, even in subtle or ambiguous cases. Recent works by Lee and Elhamifar [lee2025error], as well as Kung et al. [kung2025changed] showed that, given prior knowledge of an activity’s execution protocol, one can use structured procedural knowledge bases (e.g., WikiHow [koupaee2018wikihow]) or LLM-generated textual descriptions to define expected pre- and post-action states and hypothetical post-conditions for potential failures. These textual representations can then be aligned with visual features to reinforce plausible state transitions while decoupling them from counterfactual ones via contrastive learning, thereby strengthening the visual grounding of procedural semantics. Complementing these generative approaches, Li et al. [li2025mistake] demonstrated a retrieval-based data augmentation strategy which allows the automatic construction of semantically attributed mistake video samples by systematically cross-matching instruction texts with video segments of disparate actions from existing large-scale corpora (e.g., pairing a “pick up hammer” instruction with a “pick up bolt” video) allowing for retrieval-based clip switching. However, such strategy fails to maintain visual consistency with the preceding action clips of the activity sample, as well as preserving scene semantics. This discontinuity disrupts the temporal coherence required to model realistic error propagation and causal dependencies in long-horizon activities.
@@ IX Conclusions
Vision-based mistake analysis in procedural activities is a rapidly evolving field with potential to improve task performance, safety, and efficiency across domains. This paper reviews existing methods, datasets, and evaluation metrics, outlining challenges and opportunities in detecting and predicting procedural and executional errors. By categorizing approaches by procedural structure, supervision, and learning strategies, we offer a structured view of the state of the art. Key challenges remain, such as the ambiguity between permissible variations and true mistakes, limited comprehensive datasets, and the need for models that reason about error dependencies and propagation. Addressing these requires innovations such as neuro-symbolic reasoning, leveraging large language models for procedural understanding, and enriching datasets with counterfactual transitions and synthetic mistakes. Future research should emphasize explainable mistake analysis, stronger cross-task generalization, and real-time predictive detection, unlocking new possibilities for human–robot collaboration, automation, and assistive technologies, and ultimately transforming how mistakes are understood and mitigated in structured activities.